OpenStudy (anonymous):

The formula for the volume of a cone is V=(1/3)pi r^2h. Find the rate of change formula if dr/dt is 3 inches per minute and h=3r when (a) r=3 inches and (b) r=9 inches

OpenStudy (anonymous):

This what I got so far... Is this right? V=(1/3) pi r^2h d/dt[V] = (1/3) pi r^2h dV/dt = (1/3) 2pi r dr/dh *h dV/dt = (1/3) 2pi(3)(3)(3)(9) My answer is: dV/dt = 162 pi

OpenStudy (anonymous):

Correction: *Find the rate of change of the volume if...

OpenStudy (campbell_st):

start by substituting 3r for h so the volume is \[V = \frac{1}{3} \times \pi \times r^2 \times 3r~~~or~~~ V = \pi r^3\] then differentiate V with respect to r dV/dr next using the change rule \[\frac{dV}{dt} = \frac{dV}{dr} \times \frac{dr}{dt}\] hope it helps

OpenStudy (campbell_st):

when you get the equation for dV/dt substitute r = 3 and evaluate then the same for r = 9

OpenStudy (campbell_st):

oops chain rule... note change rule